The perimeter of a rectangle is 84cm. The length is twice the width . Find the length and width. I know that the length is 28 and the width of 14 ( a friend has tried to explain this to me) but I can't figure out the formula.
width = x
length = 2x
perimeter=84.
Can't figure it out!

Found 2 solutions by pwac, 22:
Answer by pwac(253)

An easy way of figuring this out is to do a sketch
marking the width as x and length 2x
If you mark all 4 sides you will see
x+x+2x+2x=84
2(x)+2(2x)=84 or
2x+4x=84
hope this helps
Pete

Answer by 22(3)